A ball is thrown vertically upward with a velocity of 30m/s-1


Give its velocity after 1s.


Determine the time required for the ball to reach max height.


Find the time required to reach a height of 30m.


Solve for the maximum height the ball achieves.



1
Expert's answer
2021-03-17T16:53:17-0400

(a)

"v=v_0-gt,""v(t=1\ s)=30\ \dfrac{m}{s}-9.8\ \dfrac{m}{s^2}\cdot1\ s=20.2\ \dfrac{m}{s}."

(b)

"v=v_0-gt,""0=v_0-gt,""t=\dfrac{v_0}{g}=\dfrac{30\ \dfrac{m}{s}}{9.8\ \dfrac{m}{s^2}}=3.06\ s."

(c)

"y=v_0t-\dfrac{1}{2}gt^2,""30=30t-4.9t^2,""4.9t^2-30t+30=0."

This quadratic equation has two roots: "t_1=4.86\ s" and "t_2=1.26\ s." The correct answer is "t=1.26\ s" (because "t_2" corresponds to the case when the ball returns from its maximum height to the 30 meters height).

(d)

"y_{max}=30\ \dfrac{m}{s}\cdot3.06\ s-\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(3.06\ s)^2=45.9\ m."